Tegan is trying to decide if a coin is fair. She flips it 100 times and gets 63 heads. Please explain why it might make sense to view 63 heads as enough evidence to conclude the coin is unfair.
Answer #1

here for a fair coin ; probability of getting a head p=1/2

hence for 100 flips ; expected heads =np=100*1/2 =50

std deviation =sqrt(np(1-p))=5

as 63 heads fall more then 2 std deviation from mean value of heads which is considered to be an unusual event ; therefore having 63 heads is  enough evidence to conclude the coin is unfair.